40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Aug 18 23:59:53 2017, gautam, Update, ALS, X Arm ALS lock 
    Reply  Sat Aug 19 01:35:08 2017, ericq, Update, ALS, X Arm ALS lock ALS_comparison.pdf
    Reply  Mon Aug 21 23:38:55 2017, gautam, Update, ALS, ALS out-of-loop noise ALS_21082017.pdf
       Reply  Tue Aug 22 02:19:11 2017, gautam, Update, ALS, ALS OLTFs ALS_OLTFs.pdfALS_OLTF_Aug2017.zip
          Reply  Tue Aug 22 23:27:14 2017, rana, Update, ALS, ALS OLTFs 
Message ID: 13237     Entry time: Mon Aug 21 23:38:55 2017     In reply to: 13229     Reply to this: 13238
Author: gautam 
Type: Update 
Category: ALS 
Subject: ALS out-of-loop noise 

I worked a little bit on the Y arm ALS today. 

  • Started by locking the Y arm to IR with POY, and then ran the dither alignment script to maximize Y arm transmission.
  • Green TRY DC monitor was around 0.16, whereas I have seen ~0.45 when we were doing DRFPMI locking.
  • So I went to the Y end table and tweaked the steering mirrors a little. I was able to get GTRY to ~0.42. I think this can be tweaked a little further but I decided to push on for tonight.
  • The beat amplitude on the network analyzer in the control room is comparable to the X arm beat now.
  • Adjusted the gain of the phase tracker servos, cleared phase history.
  • Looking at the ALS beat noise with the arms locked to IR and the slow ALS temperature control loops ON (see Attachment #1), the current measurements line up quite well with the reference traces.

I am now going to measure the OLTFs of both green PDH loops to check that the overall loop gain is okay, and also check the measurement against EricQ's LISO model of the (modified) AUX green PDH servos. Results to follow.


Some weeks ago, I had moved some of the Green steering optics on the PSL table around, in order to flip some mirror mounts and try and get angles of incidence closer to ~45deg on some of the steering mirrors. As a result of this work, I can see some light on the GTRY CCD when the X green shutter is open. It is unclear if there is also some scattered light on the RFPDs. I will post pictures + a more detailed investigation of the situation on the PSL table later, there are multiple stray green beams on the PSL table which should probably be dumped.


As I was writing this elog, I saw the X green lock drop abruptly. During this time, the X arm stayed locked to the IR, and the Y arm beat on the control room network analyzer did not jump (at least not by an amount visible to the eye). Toggling the X end shutter a few times, the green TEM00 lock was re-acquired, but the beatnote has moved on the control room analyzer by ~40MHz. On Friday evening however, the X green lock held for >1 hour. Need to keep an eye on this.

Attachment 1: ALS_21082017.pdf  74 kB  | Hide | Hide all
ALS_21082017.pdf
ELOG V3.1.3-